A doctor prescribes 10 mg of cough syrup. Available is 5 mg/5 mL. How many tablespoons would you give?

A. 1/2 tablespoon

B. 1 tablespoon

C. 1 1/2 tablespoons

D. 2/3 tablespoon
Answer #1

ANSWER.

* Calculation of amount (tablespoons) of cough syrup required.

          Formula = Desired dose / Dose in hand * Quantity.

              Desired dose (D) = 10 mg.

            Dose in hand (H) = 5 mg.

                   Quanitity (Q) = 5 ml.

     As per formula

            Amount of cough syrup required = D/H * Q

                                                                 = 10/5 * 5

                                                                 = 2 * 5

                                                                 = 10 ml.

        1 tablespoon = 15 ml.

                                  So

                                          2/3 tablespoon = 2/3 * 15

                                                                    = 10 ml.

      Amount(Tablespoon) of cough syrup required = 2/3 tablespoon (D)